Two force vectors with |A| = 65 N and |B| = 76 N make an angle of 45° with each other. C is the sum of these two vectors. Calculate the magnitude of C
1
Expert's answer
2020-07-22T10:35:18-0400

The Cosine Rule says

"C=\sqrt{A^2+B^2-2AB\cos \theta}=\\
\sqrt{65^2+76^2-2\times 65\times 76\cos 45^{\circ}}=55\:\rm N"